# The New Science of Unidentified Aerospace-Undersea Phenomena (UAP)

### v1

@@ A.6 Richard Lee Hoffman
During the 60 years, Rich conducted more than 1000 investigations, conducted field investigator training sessions, and held numerous additional positions with MUFON. He was the State Director for Alabama and Mississippi, the Deputy Director of Investigations and Star Team Manager from 2009 to 2015 and later the Director of Strategic Projects, a role that sought to bring technology to the forefront. In 2017, Rich left his MUFON leadership positions and cofounded the Scientific Coalition for UAP Studies (SCU). The organization has grown exponentially to now close to 400 members worldwide. The vast majority of which are scientists, engineers, researchers, military and many other disciplines. Rich is also an advisor and supporter for UAPx, the Society for UAP Studies (SUAPS).
@@
A.7 Dr. J. Allen Hynek, Northwestern University
@@
While at Ohio State University, in 1948, Hynek had begun consulting for Project Sign at Wright-Paterson Air Force Base in Dayton, Ohio to investigate reports of UFOs. Hynek was initially skeptical of reports on UFOs, and stated that ‘‘the whole subject seems utterly ridiculous.’’ Hynek continued working for the Air Force as Project Sign became Project Grudge, and later remained a scientific consultant for Project Blue Book while he was at Northwestern University.
@@
A.8 Dr. Bruce Maccabee, US Naval Surface Warfare Center
@@
Dr. Maccabee’s interest in UFOs began in the 1960s. He authored several books on the subject [384][385][110] and has written dozens of articles on specific UFO sighting reports. As an optical scientist, his analysis of the luminosity of UFOs [57], as well as his analysis of two UFO photos taken in McMinnville, Oregon [386] were invaluable . His UFO research and investigations also included the Kenneth Arnold sighting of 1947 [387], the F4 engagements of a UFO over Tehran in 1976 [388], the Kaikōura New Zealand videos and sightings of December 1978 [389, 390], the Japan Airlines (JAL1628) sighting and radar data of 1986 [391, pp. 17–18], and the Phoenix Lights sightings of March 1997 [392].
A.9 Dr. James McDonald, University of Arizona
@@
McDonald wrote detailed criticisms of the Condon report [398], and was especially disturbed by the fact that Condon, in his introduction, stated that all UFOs could be explained as hoaxes or misidentified prosaic objects, despite the fact that the report has classified 30% of the cases investigated as ‘‘unknown’’. In a letter to Condon, dated June 6, 1969, McDonald wrote: [399, Appendix 16-B, p.574]
@@
A.10 Dr. Donald Menzel, Harvard College Observatory
@@
His colleague at Harvard, Dr. Dorrit Hoffleit, recalls that one of his first actions was asking his secretary to destroy a third of the astronomical photographic plates (sight unseen) [403]. This resulted in their permanent loss from the record, which was referred to as the ‘‘Menzel Gap’’. It is important to note that the Menzel Gap adversely affects the modern efforts by Dr. Beatriz Villaroel and her VASCO Project (see 4.1.6) to identify astronomical transients and possibly connect some of them with UAP [282].
Menzel was a prominent skeptic concerning UFOs. He authored or co-authored three popular books debunking UFOs: Flying Saucers - Myth - Truth - History [404], The World of Flying Saucers [405], and The UFO Enigma [406], all of which argued that UFOs are misidentifications of prosaic phenomena, along with several papers along with papers [407, 408, 409, 410].
A.11 Dr. Richard Niemtzow, Brooks Air Force Base
@@
A.12 Dr. Claude Poher, CNES
@@
A.13 Dr. Harley Rutledge, Southeast Missouri State Univ.
@@
A.14 Dr Peter Sturrock, Stanford University
@@
A.15 Dr. Michael Swords, Western Michigan University
@@
Professor Swords is a long term board member of the J. Allen Hynek Center for UFO Studies, and he edited the Journal of UFO Studies for six years. With over 40 years of experience studying the UFO subject, Professor Swords has written extensively on the two most famous examinations of the subject: The Robertson Panel and the Condon Committee. His writings are referenced in [119], citeSwords+etal:2012, and [420].
A.16 Dr. Jacques F. Vallée, Documatica Research

### v2

@@ A.7 Richard Lee Hoffman
During the 60 years, Rich conducted more than 1000 investigations, conducted field investigator training sessions, and held numerous additional positions with MUFON. He was the State Director for Alabama and Mississippi, the Deputy Director of Investigations and Star Team Manager from 2009 to 2015 and later the Director of Strategic Projects, a role that sought to bring technology to the forefront. In 2017, Rich left his MUFON leadership positions and cofounded the Scientific Coalition for UAP Studies (SCU). The organization has grown rapidly, now close to 400 members worldwide. The vast majority are scientists, engineers, researchers, military, and many other disciplines. Rich is also an advisor and supporter for UAPx and the Society for UAP Studies (SUAPS).
@@
A.8 Dr. J. Allen Hynek, Northwestern University
@@
While at Ohio State University, in 1948, Hynek had begun consulting for Project Sign at Wright-Patterson Air Force Base in Dayton, Ohio to investigate reports of UFOs. Hynek was initially skeptical of reports on UFOs, and stated that ‘‘the whole subject seems utterly ridiculous.’’ Hynek continued working for the Air Force as Project Sign became Project Grudge, and later remained a scientific consultant for Project Blue Book while he was at Northwestern University.
@@
A.9 Dr. Bruce Maccabee, US Naval Surface Warfare Center
@@
Dr. Maccabee’s interest in UFOs began in the 1960s. He authored several books on the subject [449][450][112] and has written dozens of articles on specific UFO sighting reports. As an optical scientist, his analysis of the luminosity of UFOs [57], as well as his analysis of two UFO photos taken in McMinnville, Oregon [451] were invaluable . His UFO research and investigations also included the Kenneth Arnold sighting of 1947 [452], the F4 engagements of a UFO over Tehran in 1976 [453], the Kaikōura New Zealand videos and sightings of December 1978 [454, 455], the Japan Airlines (JAL1628) sighting and radar data of 1986 [456, pp. 17–18], and the Phoenix Lights sightings of March 1997 [457].
A.10 Dr. James McDonald, University of Arizona
@@
McDonald wrote detailed criticisms of the Condon report [463], and was especially disturbed by the fact that Condon, in his introduction, stated that all UFOs could be explained as hoaxes or misidentified prosaic objects, despite the fact that the report has classified 30% of the cases investigated as ‘‘unknown’’. In a letter to Condon, dated June 6, 1969, McDonald wrote: [464, Appendix 16-B, p.574]
@@
A.11 Dr. Donald Menzel, Harvard College Observatory
@@
His colleague at Harvard, Dr. Dorrit Hoffleit, recalls that one of his first actions was asking his secretary to destroy a third of the astronomical photographic plates (sight unseen) [468]. This resulted in their permanent loss from the record, which was referred to as the ‘‘Menzel Gap’’. It is important to note that the Menzel Gap adversely affects the modern efforts by Dr. Beatriz Villaroel and her VASCO Project (see 7.1.6) to identify astronomical transients and possibly connect some of them with UAP [339].
Menzel was a prominent skeptic concerning UFOs. He authored or co-authored three popular books debunking UFOs: Flying Saucers - Myth - Truth - History [469], The World of Flying Saucers [470], and The UFO Enigma [471], all of which argued that UFOs are misidentifications of prosaic phenomena, along with several relevant papers [472, 473, 474, 475].
A.12 Dr. Richard Niemtzow, Brooks Air Force Base
@@
A.13 Dr. Claude Poher, CNES
@@
A.14 Dr. Harley Rutledge, Southeast Missouri State Univ.
@@
A.15 Dr. Peter Sturrock, Stanford University
@@
A.16 Dr. Michael Swords, Western Michigan University
@@
Professor Swords is a long term board member of the J. Allen Hynek Center for UFO Studies, and he edited the Journal of UFO Studies for six years. With over 40 years of experience studying the UFO subject, Professor Swords has written extensively on the two most famous examinations of the subject: The Robertson Panel and the Condon Committee. His writings are referenced in [121], [67], and [485].
A.17 Dr. Jacques F. Vallée, Documatica Research